hi how would i go about solving 9^(x-5)=81^(x-7) for logarithms i know that the 81 would turn to 9 to have a common base but im not sure what else to do with the powers?
Answer by mananth(16949)

9^(x-5)=81^(x-7)
9^(x-5)=9^2(x-7)
since the bases are same
x-5 = 2(x-7)
x-5=2x-14
2x-x=14-5
x=9
...
m.ananth@hotmail.ca